# Autodeleveraging: Impossibilities and Optimization

### v1

@@ 9.1 Experimental Setup
Dataset.
We utilize the canonical reconstruction of the event [HyperMultiAssetedADL], covering a 12-minute window (21:16–21:27 UTC) with \$2.1B in liquidations across 161 assets. The replay comprises $T=1108$ ADL shocks (998 with $D_{t}>0$). Each shock records the realized deficit $D_{t}$ (total negative equity of the losers) and the account-level winner capacities $w_{t}$. Summing $D_{t}$ across the positive-deficit shocks yields \$304.5M of shortfall that must be socialized, whereas the feasible haircut mass $\sum_{i}w_{t,i}$ totals only \$95.8M, so at least \$208.6M of losses remain uncovered irrespective of policy. These pathwise statistics match the appendix methodology and provide the baseline against which we evaluate each controller. The dataset provides the exact sequence of exogenous price shocks and trader positions, allowing for a counterfactual analysis of ADL responses holding market conditions fixed.
Policies.
We compare the static benchmark $\pi_{\text{queue}}$ against the dynamic controllers defined in Appendix H.3:
  • Static Priority ($\pi_{\text{queue}},\pi_{\text{smart}}$): The standard greedy mechanism (Queue) and its deficit-capped variant (Smart Queue). $\pi_{\text{smart}}$ enforces $H_{t}\leq D_{t}$ but retains the greedy priority ranking.
  • Dynamic Severity ($\pi_{\text{exp}},\pi_{\text{md}}$): The Exponential Backoff and Mirror Descent controllers that optimize the scalar severity $\theta_{t}$ while allocating haircuts pro-rata.
  • Vector Optimization ($\pi_{\text{vec}}$): The full allocation policy minimizing the composite loss $L_{t}$ via high-dimensional mirror descent.
@@
We construct a revenue proxy that estimates the expected fee and revenue loss for the exchange given a set of ADL haircuts. At a high level, we model how much churn traders experience post haircut, akin to other customer retention models [Acerbi2002, Bolton1998Duration, Schmittlein1987Customers]. Given the churn, we model a fee proxy that estimate the loss in fees per unit of notional size that churns. Formally, define the churn probability of winner $i$ as the probability that the account exits the venue after receiving a haircut in round $t$. Empirically, this quantity controls how much future revenue the exchange forfeits. We model churn with the exponential hazard $p_{t,i}=1-\exp(-\beta h_{t,i}/w_{t,i})$, which keeps responses linear for small haircuts yet saturates toward one for severe cuts, matching the account-level attrition we measure in the Hyperliquid replay. This mirrors the survival-style retention curves widely used in both classic and modern marketing-science churn models [Bolton1998Duration, Schmittlein1987Customers, AscarzaHardie2013UsageChurn, LemmensGupta2020Churn, Ascarza2018RetentionFutility]. Full details of the calculation and the fee proxy are in Appendix H.2.
@@ 9.2 Results
Overshoot and Efficiency.
Consistent with the theoretical prediction of excessive severity in static regimes (Section 5), $\pi_{\text{queue}}$ amassed a cumulative overshoot of \$630M (Figure 3). While $\pi_{\text{smart}}$ eliminates this technical inefficiency by capping haircuts at the deficit, it does so by concentrating losses on the largest winners. This triggers the adverse selection feedback loop: $\pi_{\text{smart}}$ induces high churn among liquidity providers, paradoxically resulting in a larger residual deficit than the dynamic policies (Figure 4).
Figure 3: ADL overshoot ($H_{t}-D_{t}$). Dashed: per-shock overshoot. Solid: cumulative overshoot. The static policy $\pi_{\text{queue}}$ accumulates $\approx\mathdollar 630$M of excess haircutting. Dynamic policies ($\pi_{\text{md}},\pi_{\text{vec}}$) keep overshoot negligible by construction, while $\pi_{\text{exp}}$ maintains it within a bounded constant factor.
The ADL Trilemma.
We analyze the ADL trilemma (Proposition 2.1 empirically by looking at the components of regret for each policy. These components represent the shortfall, fairness, and revenue loss realized by using each ADL policy. Figure 7 anchors the “solvency-at-any-cost” corner of the trilemma: $\pi_{\text{queue}}$ slashes residual deficits (i.e., higher solvency) yet incurs runaway overshoot and fairness penalties (i.e., lower fairness and revenue). Figure 8 shows that capping overshoot via $\pi_{\text{smart}}$ still leaves concentrated losses (i.e., low fairness) and elevated residual deficits once churn feedback kicks in (i.e., lower long term revenue). Figures 9 and 10 illustrate how dynamic mirror-descent controllers match the queue’s residual deficits without its welfare cost, i.e., they have better revenue preserving properties than Queue while preserving approximately the same amount of solvency. On the other hand, Figure 11 demonstrates that (static) exponential backoff strategies can be competitive with dynamic policies if the deficit is sufficiently bounded. By jointly optimizing allocation and severity, Figure 12 shows $\pi_{\text{vec}}$ delivering the fairest loss sharing and the steepest long term revenue recovery.
Figure 4: Outstanding negative equity (cumulative residual deficits). Lower is better. Despite its aggression, $\pi_{\text{queue}}$ fails to clear the deficit effectively due to feedback loops. $\pi_{\text{exp}}$ achieves competitive solvency without the welfare cost.
Figure 5: Largest winner haircut (USD). $\pi_{\text{queue}}$ places a massive, spikey burden on single participants. Dynamic policies ($\pi_{\text{md}},\pi_{\text{exp}}$) distribute the load, respecting fairness constraints.
Figure 6: PTSR (top) and PMR (bottom) on log scale. $\pi_{\text{queue}}$ exhibits ratios $10^{6}$, indicating extreme moral hazard. Adaptive policies remain within the theoretical corridor ($\approx 1$$10$).
@@
Figure 7 makes clear that overshoot (orange) and fairness (red) dominate the queue baseline. Figure 8 cuts the overshoot bar almost entirely but replaces it with a larger residual (blue) mass as liquidity flees. Mirror-descent controllers in Figures 9 and 10 shrink every component simultaneously by aligning severity with observed deficits, with the warm-started variant tightening the residual tail. Figure 11 trades a small overshoot allowance for even lower fairness cost, while Figure 12 compresses all three components and traces the Pareto frontier across solvency, fairness, and revenue.
Figure 7: Regret decomposition for the production Queue baseline $\pi_{\text{queue}}$. Overshoot (orange) and fairness (red) penalties dominate, illustrating static inefficiency.
Figure 8: Regret decomposition for the Smart Queue $\pi_{\text{smart}}$. Capping overshoot helps, but churn-driven residual deficits and fairness costs still dominate regret.
Figure 9: Regret decomposition for Mirror Descent $\pi_{\text{md}}$. Adaptive severity minimizes overshoot while balancing fairness and residual losses.
Figure 10: Regret decomposition for the Dynamic2 controller $\pi_{\text{dyn2}}$. Warm-started mirror descent keeps overshoot tame while moderating fairness penalties.
Figure 11: Regret decomposition for Exponential Backoff $\pi_{\text{exp}}$. Feedback-tuned severity contains overshoot without sacrificing solvency or fairness.
Figure 12: Regret decomposition for Vector Mirror $\pi_{\text{vec}}$. Joint severity and allocation updates compress every component, tracing the Pareto frontier.
@@ 9.3 Interpretation
In aggregate, the empirical evidence of this section strongly suggests that the standard Queue mechanism operates in the unbounded PoA regime defined in Proposition 8.3. It fails to balance the competing objectives of a perpetuals exchange, sacrificing trader welfare and future revenue for marginal (and often negative) gains in immediate solvency. The decoupling of severity ($\theta_{t}$) from allocation ($h_{t}$) via policies like Exponential Backoff ($\pi_{\text{exp}}$) offers a strictly dominant strategy, improving fairness and revenue retention without compromising system stability.
Based on these results, we offer three concrete prescriptions for Hyperliquid and similar venues:
@@
    Decouple Severity from Allocation. Current mechanisms conflate the decision of how much to haircut with who to haircut. We recommend adopting a scalar feedback controller (e.g., $\pi_{\text{exp}}$ or $\pi_{\text{md}}$) to determine the aggregate severity $\theta_{t}$ dynamically based on the observed deficit $D_{t}$.
@@
    Adopt Fairer Allocations. The greedy ranking used by $\pi_{\text{queue}}$ maximizes the fairness penalty and adverse selection. Shifting to a Levered Pro-Rata (LPR) rule (or the optimized vector $h_{t}$ from $\pi_{\text{vec}}$) ensures that loss socialization is proportional to risk and capacity, maintaining $\text{PTSR}\approx 1$.
@@
    Implement a Hybrid Controller. We recommend a hybrid approach: use LPR (Eq. 21) for allocation to satisfy fairness constraints, paired with the $\pi_{\text{exp}}$ controller for severity to manage the deficit. This combination provides the robustness of dynamic severity with the transparency and fairness of pro-rata allocation, minimizing the long-term revenue impact of ADL events.
@@ 10 Conclusion and Future Work
In this paper, we provide the first formalism for ADL mechanisms that allows us to compare the performance of different policies. We first that a negative result: it is impossible for a perpetuals to optimize for exchange solvency, long term revenue, and fairness to traders. We then try to get around the negative result by analyzing strategies that optimize each of these components individually. Our results demonstrate that the heuristic strategy employed by Hyperliquid, Binance, and others that stems from 2015 is actually suboptimal in all dimensions. We empirically validate our results on Hyperliquid data from October 10, 2025 and demonstrate the superiority of dynamic and even other static algorithms to the live policy.
@@
ADL policies also likely change in their efficacy if they were able to be executed in a fully privately manner. In this paper, we ignored modeling execution costs for exchanges upon liquidation and position closure. The modeling of such execution costs is nuanced and likely leads to incorrect conclusions about ADL if not carefully handled [^12] See, e.g. Storm2025ADLThread1, Storm2025ADLThread2 for examples of overly naive analyses that conflate execution costs with ADL mechanism design.. When private execution is possible, the impact of these costs is muted as follower strategies are not able to react to ADL shocks.
@@
Finally, we note that the results in this paper focused on a single margin model, where the posted collateral was the numéraire. Many exchanges, including Hyperliquid, have cross-margin ADL support. This changes the modeling of execution costs and forces us to explicitly model market impact on cash / collateral balances.

### v3

@@ 9.1 Experimental Setup
Regret Framework.
We evaluate ADL policies using two complementary regret components:
Definition 9.1 (Fairness Regret).
The fairness regret measures the maximum deviation from pro-rata allocation:
$\mathcal{R}_{t}^{\mathrm{fair}}(\pi):=\max_{i\in W_{t}}\frac{h_{\pi,i}}{e_{i}}-\frac{h_{\mathrm{pr},i}}{e_{i}}$
where $h_{\pi,i}$ is the haircut under policy $\pi$ and $e_{i}$ is user $i$’s equity.
Definition 9.2 (Overshoot Regret).
The overshoot regret measures excess budget beyond the needed amount:
$\mathcal{R}_{t}^{\mathrm{overshoot}}(\pi):=B_{t}^{\pi}-B_{t}^{\mathrm{needed}}$
Total regret is the sum of these components. Policies with low overshoot but highly concentrated losses, and policies with diffuse losses but systematic overshoot, are both suboptimal under this decomposition.
Dataset and wave partition.
We utilize the canonical public-data reconstruction of the event [42], covering a 12-minute window (21:16–21:27 UTC) with \$2,103,111,431 in liquidations across roughly 160 assets. We partition ADL fills into global time waves by gap clustering: a new wave begins when the inter-fill gap exceeds 5 seconds, yielding $T=16$ waves for this window.
Deficit and haircut capacity.
For each wave $t$, the realized loser deficit is
$D_{t}=\sum_{j\in\mathrm{losers}(t)}(-e_{t}(j))_{+},$
computed from loser-side liquidation equity fields (using the canonical table’s liquidated_total_equity, with a per-liquidated-user minimum within a wave to avoid double counting). Across the event, the realized deficit sums to $\sum_{t}D_{t}\approx\mathdollar 100.1$M. Haircuts apply to profits (PNL), not protected cash/collateral. For benchmark allocations we use a per-user capacity proxy $c_{u}:=\min(U(u),E(u))$, where $U(u)$ is positive unrealized PNL and $E(u)$ is positive equity, both taken from the canonical REALTIME table within the wave window.
Needed budget and production wealth removed.
We compute $B_{t}^{\mathrm{needed}}$ from mark–execution gaps in the raw fill stream and $H_{t}^{\mathrm{prod}}(\Delta)$ via two-pass replay as described below. At $\Delta=0$, totals satisfy $\sum_{t}B_{t}^{\mathrm{needed}}\approx\mathdollar 15.1$M and $\sum_{t}H_{t}^{\mathrm{prod}}\approx\mathdollar 60.1$M, hence $O(0)\approx\mathdollar 45.0$M.
Regime check .
To assess whether the October 10 episode plausibly lies in the structural-deficit regime ($\mu_{-}\gg\mu_{\Phi}$), we run an order-of-magnitude check. The realized event deficit is $\sum_{t}D_{t}\approx\mathdollar 100.1$M. Using the venue fee schedule and event turnover, a conservative upper bound for sustainable fee diversion is approximately 2–5% of gross fees; with gross fees in the \$1–5M range over the window, this implies $\mu_{\Phi}$ on the order of \$20k–\$250k. The realized deficit is therefore roughly two orders of magnitude larger, consistent with $\mu_{-}\gg\mu_{\Phi}$ in this episode. This is a plausibility check, not identification: the formal trilemma statement still relies on Assumption J.3.
Benchmark allocations.
We compare production to transparent benchmark allocations that target $B_{t}^{\mathrm{needed}}$ on each wave under the same profits-only capacity constraint:
  • Wealth pro-rata (continuous): a capped pro-rata allocation in wealth-space USD.
  • Vector mirror descent (vector-md): a projection-based allocator that matches the budget using a single projection step onto $\{x\in[0,1]^{n}:c^{\top}x=B\}$.
  • Contract pro-rata: standard exchange-style integer allocation proportional to position size, used by Binance, Bybit, and similar venues.
  • Min-max ILP: MIP solver minimizing the maximum haircut percentage $\max_{i}h_{i}/e_{i}$ subject to budget and capacity constraints.
These benchmarks are used to quantify the mechanically necessary overshoot under discreteness and capacity constraints; they are not claims about production implementability.
Integer rounding and budget shortfall.
A key difference between continuous and integer-contract allocation is rounding error. In contract space, standard pro-rata rounding (downward rounding of fractional contracts) can induce substantial under-allocation relative to $B_{t}^{\mathrm{needed}}$. In our reconstruction, contract pro-rata exhibits per-wave shortfalls up to \$371,000 (wave 1) and cumulative under-allocation of approximately \$1.5M across the episode. Thus, standard exchange-style integer allocation does not reliably hit the target budget without additional adjustment. The min-max ILP benchmark enforces budget matching while minimizing the maximum haircut percentage, achieving near-exact targeting with only small positive overshoot (cumulative \$50,000). This shows that tight budget control is feasible in integer contract space, but not via simple rounding alone.
@@
We construct a revenue proxy to estimate expected fee loss induced by ADL haircuts. The proxy maps haircut intensity to post-event account churn, consistent with retention modeling in the economics and marketing literature [2, 26, 102]. Formally, let the churn probability for winner $i$ in wave $t$ be the probability that the account exits after receiving haircut $h_{t,i}$. We model this probability with an exponential hazard:
$p_{t,i}=1-\exp(-\beta h_{t,i}/w_{t,i}).$
This specification is approximately linear for small haircuts and saturates toward one for severe haircuts, matching the account-level attrition patterns observed in the replay. It is also consistent with survival-style retention curves used in both classical and modern churn studies [26, 102, 14, 81, 15]. Full details of the calculation and the fee proxy are in Appendix I.3.
Overshoot measurement methodology.
To measure overshoot, we compare production ADL (executed in contract space) with wealth-space comparator policies. This requires a two-pass replay on the same realized price path.
Two-pass replay. For each ADL wave $t$, we define two equity outcomes for each position $p$:
  • $e_{t}^{\mathrm{ADL}}(p)$: the realized equity after production ADL executes in wave $t$ (observed outcome).
  • $e_{t}^{\mathrm{no\text{-}ADL}}(p)$: a counterfactual equity if ADL were disabled in wave $t$, holding the realized price path fixed (reconstructed outcome).
The induced haircut from production ADL is then:
$h_{t}^{\mathrm{prod}}(p)=\begin{cases}\dfrac{e_{t}^{\mathrm{no\text{-}ADL}}(p)-e_{t}^{\mathrm{ADL}}(p)}{e_{t}(p)},&e_{t}(p)>0,\\[6.0pt] 0,&\text{otherwise},\end{cases}$
where $e_{t}(p)$ is the pre-ADL equity (which equals $e_{t}^{\mathrm{no\text{-}ADL}}(p)$ by construction of the counterfactual). Total wealth removed from winners in wave $t$ is:
$H_{t}^{\mathrm{prod}}=\sum_{p\in\mathcal{W}_{t}}\bigl(e_{t}^{\mathrm{no\text{-}ADL}}(p)-e_{t}^{\mathrm{ADL}}(p)\bigr)_{+}.$
The no-ADL counterfactual and identification assumption. The counterfactual $e_{t}^{\mathrm{no\text{-}ADL}}(p)$ is not directly observed because production ADL executed. We recover it by replaying the realized price path with ADL disabled, under the short-horizon assumption that positions remain open and equity evolves with marks over the execution window. This assumption is necessary because trader actions in the unobserved world (manual closes, leverage changes, transfers) are not observed. Given that ADL waves occur on a seconds-scale during high volatility, this assumption is appropriate for measuring immediate wealth transfer effects.
Needed budget and production overshoot vs needed. For each ADL fill $k$ with liquidation mark $p_{k}^{\mathrm{mark}}$, execution price $p_{k}^{\mathrm{exec}}$, and size $q_{k}$, define the instantaneous bankruptcy-gap proxy
$\mathrm{needed}_{k}:=\lvert p_{k}^{\mathrm{mark}}-p_{k}^{\mathrm{exec}}\rvert\cdot\lvert q_{k}\rvert$.
We aggregate within each wave $t$:
$B_{t}^{\mathrm{needed}}:=\sum_{k\in\text{fills in wave }t}\mathrm{needed}_{k}.$
We define the production overshoot vs needed:
$O_{t}(\Delta):=H_{t}^{\mathrm{prod}}(\Delta)-B_{t}^{\mathrm{needed}},\qquad O(\Delta):=\sum_{t}O_{t}(\Delta).$
This compares realized wealth removed from winners to the minimal instantaneous transfer implied by closing across the mark–bankruptcy gap.
Evaluation horizon $\Delta$. The quantity $B_{t}^{\mathrm{needed}}$ is computed from realized ADL fills and is independent of $\Delta$. By contrast, $\Delta$ changes the evaluation time for winner equities in $H_{t}^{\mathrm{prod}}(\Delta)$: we evaluate $e^{\mathrm{ADL}}$ and $e^{\mathrm{no\text{-}ADL}}$ at $t_{\mathrm{eval}}=t_{\mathrm{end}}+\Delta$ while holding the realized post-wave path fixed. This isolates a short-horizon opportunity-cost channel. Over $\Delta\in\{0,500,1000,2000,5000\}$ms, we obtain $O(0)\approx\mathdollar 45.0$M and a range of approximately \$45.0M–\$51.7M.
@@ 9.2 Results
Headline numbers (two numeraires).
Figure 3 reports two headline numeraires: the stylized queue overshoot diagnostic in equity dollars (wealth space) and the production overshoot vs needed metric in PNL dollars (two-pass replay at $\Delta=0$). The figure also reports the simple mapping obtained by dividing the equity-dollar queue diagnostic by the trimmed-mean winner overcollateralization ratio (equity/PNL).
[…]
Overshoot vs horizon (opportunity-cost sensitivity).
Figure 4 plots $O(\Delta)$ from the horizon sweep. As $\Delta$ increases, $H_{t}^{\mathrm{prod}}(\Delta)$ may rise because positions are marked later along the realized path, revealing the short-horizon opportunity-cost channel.
Figure 4: Total production overshoot vs needed $O(\Delta)$ as a function of evaluation horizon $\Delta$. The overshoot increases from \$45.0M to \$51.7M as $\Delta$ grows, reflecting opportunity-cost channels: positions marked later along the realized price path capture additional value that would have accrued to winners.
Production vs benchmark allocations (per wave).
To quantify excess relative to transparent alternatives, we compare production with benchmark allocations that target $B_{t}^{\mathrm{needed}}$ on each wave. Benchmarks use a PNL-only haircut-capacity proxy constructed from the canonical REALTIME table (Appendix I). Figure 5 shows per-wave budgets, overshoot vs needed, and a concentration diagnostic.
Figure 5: Production vs benchmark allocations per wave. The “needed” budget $B_{t}^{\mathrm{needed}}$ (bankruptcy-gap proxy) is compared against production haircuts $H_{t}^{\mathrm{prod}}$ and four benchmark policies: wealth pro-rata, vector mirror descent, contract pro-rata, and min-max ILP. Benchmarks achieve near-exact budget matching while production consistently overshoots, empirically validating the trilemma’s prediction that queue-based allocation sacrifices efficiency (Proposition 2.1).
Cumulative overshoot vs needed.
Figure 6 shows the aggregate effect. Benchmarks that target $B_{t}^{\mathrm{needed}}$ maintain near-zero cumulative overshoot vs needed (up to integer rounding and capacity caps), while production accumulates substantial excess.
Figure 6: Cumulative overshoot vs needed over the 16 waves. Production accumulates \$45.0M in excess haircuts while benchmark policies maintain near-zero cumulative overshoot (bounded only by integer rounding and capacity constraints). This demonstrates that large overshoot is not mechanically inevitable—it is a policy choice.
@@
We decompose total regret into overshoot regret and fairness regret, as defined in the Regret Framework. Figure 7 shows cumulative regret trajectories over the 16 waves for production versus benchmark allocations. Figures 810 provide detailed breakdowns of each regret component separately.
[…]
Figure 8: Cumulative overshoot regret by policy. Production accumulates \$45.0M in overshoot (excess haircuts beyond needed), while benchmark policies achieve near-zero overshoot. This confirms that large overshoot is a policy choice, not a mechanical necessity.
Figure 9: Cumulative fairness regret by policy. Production’s regret is dominated by fairness violations: the worst-hit winner receives 75.45% haircuts on average, compared to 11.53% for contract pro-rata and 6.21% for min-max ILP. Vector-MD achieves the lowest fairness regret (0.71%).
Figure 10: Total cumulative regret (overshoot + fairness) by policy. The dominant contributor to production’s total regret is fairness violations, not overshoot. Benchmark policies achieve bounded total regret through proportional allocation.
@@ 9.3 Interpretation
Taken together, the reconstruction indicates that production ADL removes materially more winner wealth than the instantaneous bankruptcy-gap proxy $B_{t}^{\mathrm{needed}}$, even at the conservative horizon $\Delta=0$. At the same time, transparent benchmarks (wealth-space pro-rata and solver-free contract-space approximations) can track $B_{t}^{\mathrm{needed}}$ closely and keep overshoot vs needed near zero. This difference is descriptive rather than normative: it measures how a contract-space production policy maps into wealth transfers under partial observability. Under this measurement, production exhibits both high overshoot regret (\$45.0M vs needed) and high fairness regret (loss concentration on top winners), while benchmark allocations maintain near-zero overshoot regret with bounded fairness regret.
Based on these results, we offer three concrete prescriptions for venues that use ADL:
@@
    Make the numéraire explicit. When reporting “haircut dollars,” state whether the unit is equity dollars (wealth space) or PNL dollars (haircut-capacity space), and report both when possible.
@@
    Separate eligibility/capacity from allocation. Enforce “profits-only” capacity constraints (haircuts apply to PNL endowment, not cash) and then choose an allocation rule conditional on that feasible set.
@@
    Publish a budget target. A publicly specified per-wave target such as $B_{t}^{\mathrm{needed}}$ makes it possible to evaluate (and improve) execution policies; benchmark allocations show that large overshoot is not mechanically inevitable.
@@ 10 Conclusion and Future Work
In this paper, we provide the first formalism for ADL mechanisms that allows us to compare the performance of different policies. We first prove a negative result: it is impossible for a perpetual venue to simultaneously optimize for exchange solvency, long-term revenue, and fairness to traders. We then try to get around the negative result by analyzing strategies that optimize each of these components individually. Our results demonstrate that the heuristic strategy employed by Hyperliquid, Binance, and others that stems from 2015 is actually suboptimal in all dimensions. We empirically validate the measurement framework on Hyperliquid data from October 10, 2025: the two-pass reconstruction shows a production profit-haircut overshoot vs the instantaneous bankruptcy-gap proxy $B_{t}^{\mathrm{needed}}$ on the order of tens of millions of dollars, while transparent benchmark allocations can target $B_{t}^{\mathrm{needed}}$ closely and thereby keep overshoot vs needed near zero.
@@
ADL policies also likely change in their efficacy if they were able to be executed in a fully privately manner. In this paper, we ignored modeling execution costs for exchanges upon liquidation and position closure. The modeling of such execution costs is nuanced and likely leads to incorrect conclusions about ADL if not carefully handled [^14] See, e.g. Storm [109, 110] for examples of overly naive analyses that conflate execution costs with ADL mechanism design.. When private execution is possible, the impact of these costs is muted as follower strategies are not able to react to ADL shocks.
@@
Finally, we note that the results in this paper focused on a single margin model, where the posted collateral was the numéraire. Many exchanges, including Hyperliquid (which recently launched portfolio margin [70]), have cross-margin ADL support. This changes the modeling of execution costs and forces us to explicitly model market impact on cash and/or collateral balances.